# Mildly Super-Eddington Accretion Onto Slowly-Spinning Black Holes Explains the X-Ray Weakness of the Little Red Dots

### v1

@@ Abstract
JWST has revealed a population of low-luminosity AGN at $z>4$ in compact, red hosts (the “Little Red Dots”, or LRDs), which are largely undetected in X-rays. We investigate this phenomenon using GRRMHD simulations of super-Eddington accretion onto a SMBH with $M_{\bullet}=10^{7}\,{\rm M_{\odot}}$ at $z\sim 6$, representing the median population; the SEDs that we obtain are intrinsically X-ray weak. The highest levels of X-ray weakness occur in SMBHs accreting at mildly super-Eddington rates ($41.4<\,{f_{\rm Edd}}<$ 4) with zero spin, viewed at angles $>30^{\circ}$ from the pole. X-ray bolometric corrections in the observed $2-10$ keV band reach $\sim 10^{4}$ at $z=6$, $\sim 5$ times higher than the highest constraint from X-ray stacking. Most SEDs exhibit $\,{\alpha_{\rm ox}}$ values outside standard ranges, with X-ray weakness increasing with optical-UV luminosity; they are also extraordinarily steep and soft in the X-rays (median photon index $\Gamma=3.1$, mode of $\Gamma=4.4$). SEDs strong in the X-rays have harder spectra with a high-energy bump when viewed near the hot ($>10^{8}$ K) and highly-relativistic jet, whereas X-ray weak SEDs lack this feature. Viewing a SMBH within $10^{\circ}$ of its pole, where beaming enhances the X-ray emission, has a $\sim 1.5\%$ probability, matching the LRD X-ray detection rate. Next-generation observatories like AXIS will detect X-ray weak LRDs at $z\sim 6$ from any viewing angle. Although many SMBHs in the LRDs are already estimated to accrete at super-Eddington rates, our model explains $50\%$ of their population by requiring that their masses are overestimated by a mere factor of $\sim 3$. In summary, we suggest that LRDs host slowly spinning SMBHs accreting at mildly super-Eddington rates, with large covering factors and broad emission lines enhanced by strong winds, providing a self-consistent explanation for their X-ray weakness and complementing other models.
@@ 1 Introduction
A newly discovered, enigmatic population of galaxies in the high-$z$ Universe is challenging this fact: the “Little Red Dots” (LRDs). LRDs are a population of high-$z$ sources detected in many JWST deep fields (Labbé et al., 2023; Kocevski et al., 2023; Harikane et al., 2023; Maiolino et al., 2023; Matthee et al., 2024), with the bulk of their distribution in the redshift range $84<z<$ 8, and a median of $z\sim 6$ (Kocevski et al., 2024; Akins et al., 2024).
@@
A majority of these sources do show telltale signs of the presence of an active SMBH. Greene et al. (2024) found that $\sim 60\%$ of the LRDs investigated display definitive evidence of broad emission lines, with $\rm H\alpha$ FWHMs of $>2000\,\rm km\,s^{-1}$: a signature of Type-1 AGN, for which an estimate of the mass can be derived (Greene & Ho, 2005). The absence of broad forbidden lines (e.g., the [OIII]5007) suggests that such emission cannot be associated with outflows (Maiolino et al., 2024). If the LRDs contain a central SMBH, their SEDs are assembled by a dust-obscured AGN and a young, blue galaxy (Kocevski et al., 2023).
@@
To conclude, King (2024) suggested that LRDs may be highly-beamed sources, fed at or above the Eddington rate. The resulting gas outflow velocities of $\sim 0.1-0.2c$ would make virial mass estimates difficult and lead to an overestimate. Similarly, based on semi-empirical models of the emission in different accretion regimes, Lupi et al. (2024) also suggested that the masses of the SMBHs in the LRDs may be overestimated due to accretion at largely super-Eddington rates.
@@ 2 A Primer on Accretion Regimes
The Eddington ratio $\,{f_{\rm Edd}}$ is defined as the ratio between the mass accretion rate onto the black hole and the corresponding Eddington rate:
$\,{f_{\rm Edd}}=\frac{\dot{M_{\bullet}}}{\dot{M}_{\rm Edd}}\,.$ (1)
@@
In particular, $\dot{M}_{\rm Edd}$ is the mass accretion rate at which a relativistic thin accretion disk (Novikov & Thorne, 1973) would radiate at the Eddington luminosity. $\,{f_{\rm Edd}}$ is the main parameter that defines the accretion regime and, ultimately, the characteristics of the accretion disk, jet, and the shape of the emerging spectral luminosity.
@@
For $10^{-2}\lesssim\,{f_{\rm Edd}}\lesssim 0.5$, a geometrically-thin, optically thick, radiatively efficient accretion disk is present (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973). This accretion mode is often referred to as the “standard disk” model and is the most studied for black holes spanning the entire mass range. The “bias” towards this accretion regime is likely caused by the combination of two effects: black holes accreting in this regime are sufficiently bright to be detected but, at the same time, do not require extremely large (thus, infrequent at $z=0$) reservoirs of cold, accretable gas. In the radiatively efficient regime, typical matter-to-energy conversion factors $\epsilon$ are of the order $\epsilon\sim 0.1$ (depending on the spin, Bardeen 1970; Novikov & Thorne 1973), where the emitted luminosity is $L=\epsilon\dot{M}c^{2}$.
@@
For $\,{f_{\rm Edd}}\lesssim 10^{-2}$ and $\,{f_{\rm Edd}}\gtrsim 0.5$, the disk becomes radiatively inefficient. In the lower range ($\,{f_{\rm Edd}}\lesssim 10^{-2}$), an extremely hot, two-temperature, optically-thin, advection-dominated accretion flow (ADAF) is formed, which is geometrically thick and has typical values of $\epsilon\ll 0.1$ (Narayan & Yi, 1994, 1995; Abramowicz et al., 1995; Narayan & McClintock, 2008; Yuan & Narayan, 2014). What little radiation is present in this accretion state has hardly any effect on the dynamics of the optically thin gas. Sgr A* is a classic example of a SMBH accreting in ADAF mode (Yuan et al., 2003). These systems often accumulate a strong magnetic field at small radii around the black hole to form a Magnetically Arrested Disk (MAD, Igumenshchev et al., 2003; Narayan et al., 2003), and in this state they produce relativistic jets extending to thousands of gravitational radii $r_{g}=2GM_{\bullet}/c^{2}$ if the black hole is spinning (Tchekhovskoy et al., 2011). It is likely that most black holes, at least in the local Universe, accrete in the ADAF mode and are in the MAD state (Narayan et al., 2022).
@@
For $\,{f_{\rm Edd}}\gtrsim 0.5$, the accretion disk is radiatively inefficient because of photon trapping (Begelman, 1979; Abramowicz et al., 1988). The disk is again geometrically thick, prompting the name “slim disk” for these systems. Jets are prominent in slim disks, especially if they achieve the MAD state and the black holes are spinning (Narayan et al., 2017; Curd & Narayan, 2023; Ricarte et al., 2023). The latter systems will be extremely X-ray-bright for pole-on observers because of relativistically beamed radiation from the hot disk and jet (Curd & Narayan, 2019).
@@
Radiation is critical for the dynamics and thermodynamics of both standard and slim disks (i.e., for the entire range $\,{f_{\rm Edd}}\gtrsim 10^{-2}$). There is a strong coupling between the gas and the radiation, while radiative pressure inflates the disk in the vertical direction as $\,{f_{\rm Edd}}$ approaches or exceeds unity (Abramowicz et al., 1988). Additionally, jets and winds driven only by radiation can occur in super-Eddington systems (Sądowski et al., 2014). For these reasons, a GRRMHD code, i.e., a GRMHD code with a self-consistent treatment of the radiation field, is required. The work reported in this paper uses the GRRMHD code KORAL (Sądowski et al., 2013, 2014).
@@ 3.1 GRRMHD Simulations with KORAL
The super-Eddington accretion process onto SMBHs is simulated using the GRRMHD code KORAL (Sądowski et al., 2013, 2014), which incorporates the effects of gas dynamics, magnetic fields, and radiation within a fixed gravitational field. These simulations are performed using the Kerr metric in Kerr-Schild coordinates and employ the M1 closure method (Levermore, 1984) to handle radiative transfer, along with a radiative viscosity term to address limitations inherent in the M1 scheme (Sądowski et al., 2015). The simulations include radiative processes such as synchrotron emission, free-free and bound-free (taken from Sutherland & Dopita, 1993) emission and absorption, and Compton scattering. All these simulations are performed in 3D to correctly resolve the magnetorotational instability (Balbus & Hawley, 1991).
@@ 3.2 Radiation Post-Processing with HEROIC
In the present work, post-processing involves two stages. First, HEROIC takes the time-averaged (over the final $5000t_{g}$) and axisymmetrized data from KORAL for the gas density, temperature, and four-velocity, as well as the magnetic field strength. To be conservative, it caps the gas temperature [^1] This temperature cap prevents numerical errors in isolated cells in KORAL from causing artificially high temperatures that affect the inverse Compton scattering in the output SEDs. The calculated spectra hardly change relative to not capping the temperature, but we do it for safety. at $10^{9}$ K, which is the typical coronal temperature in Seyfert galaxies (see, e.g., Akylas & Georgantopoulos 2021). Using this information and the appropriate Kerr spacetime metric, the code iteratively solves for the frequency-dependent radiation field at each grid point, including all the emission, absorption, and scattering opacities and the effect of relativistic Comptonization. At the conclusion of this stage, the code generates a source function at each position on the grid. Using this source function, in the second stage HEROIC performs general relativistic ray-tracing, again with all opacities and scattering included, to compute the observed spectra from different viewing angles.
The observed SEDs are re-normalized so that their bolometric luminosity matches the value expected from the Eddington ratio of each specific run. This procedure is standard practice to correct minor discrepancies in calculating the radiation and temperature fields performed by KORAL, using the M1 closure method. The fact that a re-normalization of a factor of $10$ was needed for the radiation field suggests that KORAL overestimated the temperature field by a factor of $10^{1/4}\approx 1.8$. A similar radiation and temperature field matching was also applied to other studies based on KORAL and HEROIC, e.g., Narayan et al. (2017).
@@ 3.3 Description of the Simulations
The GRRMHD simulations vary the Eddington ratio, ranging from mildly super-Eddington (i.e., $\,{f_{\rm Edd}}=1.4$) to strongly super-Eddington (i.e., $\,{f_{\rm Edd}}=13.4$). Three values of the dimensionless SMBH spin parameter $a$ are considered: 0, 0.68, 0.9. By definition, $a$ is required to lie in the range $0\leq|a|\leq 1$, where $a=0$ is a Schwarzschild (i.e., nonrotating) black hole, and $a=1$ is a maximally rotating black hole (although the limiting value $a=1$ is not achievable in practice, Thorne 1974).
These 12 simulations are then post-processed to obtain the emerging spectral luminosity from 8 different viewing angles, as measured by an observer at an inclination angle $i$ from the pole. The inclination angle varies between $i=10^{\circ}$ to $i=80^{\circ}$, in intervals of $10^{\circ}$. Hence, a grid of $12$ simulations $\times$ $8$ angles leads to a total of $96$ SEDs, which are used to properly investigate the physical parameter spaces relevant to our problem. A summary of the physical properties of the 12 simulations performed is presented in Table 1.

### v2

@@ Abstract
JWST has revealed a population of low-luminosity AGN at $z>4$ in compact, red hosts (the “Little Red Dots”, or LRDs), which are largely undetected in X-rays. We investigate this phenomenon using GRRMHD simulations of super-Eddington accretion onto a SMBH with $M_{\bullet}=10^{7}\,{\rm M_{\odot}}$ at $z\sim 6$, representing the median population; the SEDs that we obtain are intrinsically X-ray weak. The highest levels of X-ray weakness occur in SMBHs accreting at mildly super-Eddington rates ($41.4<\,{f_{\rm Edd}}<$ 4) with zero spin, viewed at angles $>30^{\circ}$ from the pole. X-ray bolometric corrections in the observed $2-10$ keV band reach $\sim 10^{4}$ at $z=6$, $\sim 5$ times higher than the highest constraint from X-ray stacking. Most SEDs are extraordinarily steep and soft in the X-rays (median photon index $\Gamma=3.1$, mode of $\Gamma=4.4$). SEDs strong in the X-rays have harder spectra with a high-energy bump when viewed near the hot ($>10^{8}$ K) and highly-relativistic jet, whereas X-ray weak SEDs lack this feature. Viewing a SMBH within $10^{\circ}$ of its pole, where beaming enhances the X-ray emission, has a $\sim 1.5\%$ probability, matching the LRD X-ray detection rate. Next-generation observatories like AXIS will detect X-ray weak LRDs at $z\sim 6$ from any viewing angle. Although many SMBHs in the LRDs are already estimated to accrete at super-Eddington rates, our model explains $50\%$ of their population by requiring that their masses are overestimated by a mere factor of $\sim 3$. In summary, we suggest that LRDs host slowly spinning SMBHs accreting at mildly super-Eddington rates, with large covering factors and broad emission lines enhanced by strong winds, providing a self-consistent explanation for their X-ray weakness and complementing other models.
@@ 1 Introduction
A newly discovered, enigmatic population of galaxies in the high-$z$ Universe is challenging this fact: the “Little Red Dots” (LRDs). LRDs are a population of high-$z$ sources detected in many JWST deep fields (Labbé et al., 2023; Kocevski et al., 2023; Harikane et al., 2023; Maiolino et al., 2023; Matthee et al., 2024), with the bulk of their distribution in the redshift range $84<z<$ 8, and a median of $z\sim 6$ (Kocevski et al., 2024; Akins et al., 2024; Kokorev et al., 2024).
@@
A majority of these sources do show telltale signs of the presence of an active SMBH. Greene et al. (2024) found that $\sim 60\%$ of the LRDs investigated display definitive evidence of broad emission lines, with $\rm H\alpha$ FWHMs of $>2000\,\rm km\,s^{-1}$: a signature of Type-1 AGN, for which an estimate of the mass can be derived (Greene & Ho, 2005). The absence of broad forbidden lines (e.g., the [OIII]5007) suggests that such emission cannot be associated with outflows (Maiolino et al., 2024). If the LRDs contain a central SMBH, their SEDs are assembled by a dust-obscured AGN and a young, blue galaxy (Kocevski et al., 2023). Pérez-González et al. (2024) argue that many LRDs are extremely intense and compact starburst galaxies, which produce a significant amount of dust; their global energy output is dominated by emission from OB stars and an obscured AGN.
@@
To conclude, King (2024) suggested that LRDs may be highly-beamed sources, fed at or above the Eddington rate, similar to the case of ultraluminous X-ray sources (ULXs). The resulting gas outflow velocities of $\sim 0.1-0.2c$ would make virial mass estimates difficult and lead to an overestimate. Interestingly, the Galactic X-ray binary Cyg X-3, which may be viewed as a ULX from observers located along the axis of the funnel, shows signs of X-ray weakness (Veledina et al., 2024). Additionally, based on semi-empirical models of the emission in different accretion regimes, Lupi et al. (2024) also suggested that the masses of the SMBHs in the LRDs may be overestimated due to accretion at largely super-Eddington rates.
@@ 2 A Primer on Accretion Regimes
The Eddington ratio $\,{f_{\rm Edd}}$ is defined as the ratio between the mass accretion rate $\dot{M_{\bullet}}$ onto the black hole and the corresponding Eddington rate:
$\,{f_{\rm Edd}}=\frac{\dot{M_{\bullet}}}{\dot{M}_{\rm Edd}}\,,$ (1)
@@
where $\dot{M}_{\rm Edd}$ is the mass accretion rate at which a relativistic thin accretion disk (Novikov & Thorne, 1973) would radiate at the Eddington luminosity. $\,{f_{\rm Edd}}$ is the main parameter that defines the accretion regime and, ultimately, the characteristics of the accretion disk, jet, and the shape of the emerging spectral luminosity.
@@
For $10^{-2}\lesssim\,{f_{\rm Edd}}\lesssim 1$, a geometrically-thin, optically thick, radiatively efficient accretion disk is present (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973) [^1] The inner parts of the disk are geometrically thick and somewhat radiatively inefficient already by $\,{f_{\rm Edd}}\approx 0.5$, but for simplicity we take the limit to be $\,{f_{\rm Edd}}\approx 1$.. This accretion mode is often referred to as the “standard disk” model and is the most widely studied for black holes spanning the entire mass range. The “bias” towards this accretion regime is likely caused by the combination of two effects: black holes accreting in this regime are sufficiently bright to be detected but, at the same time, do not require extremely large (thus, infrequent at $z=0$) reservoirs of cold, accretable gas. In the radiatively efficient thin disk regime, typical matter-to-energy conversion factors $\epsilon$ are of the order $\epsilon\sim 0.1$ (depending on the spin, Bardeen 1970; Novikov & Thorne 1973), where the emitted luminosity is $L=\epsilon\dot{M}c^{2}$.
The standard thin disk regime is itself sub-divided into two sub-regimes. For $\,{f_{\rm Edd}}\lesssim 0.1$, the disk is gas-pressure dominated and is thermally stable. However, for $0.1\lesssim\,{f_{\rm Edd}}\lesssim 1$, radiation pressure dominates, and the disk becomes viscously and thermally unstable (Lightman & Eardley, 1974; Shakura & Sunyaev, 1976; Jiang et al., 2013). Unfortunately, the unstable regime overlaps with the classic luminosity range of quasars, which prevents self-consistent modeling of these astrophysically important systems.
@@
For $\,{f_{\rm Edd}}\lesssim 10^{-2}$ and $\,{f_{\rm Edd}}\gtrsim 1$, the disk becomes radiatively inefficient. In the lower range ($\,{f_{\rm Edd}}\lesssim 10^{-2}$), an extremely hot, two-temperature, optically-thin, advection-dominated accretion flow (ADAF) is formed, which is geometrically thick and has typical values of $\epsilon\ll 0.1$ (Narayan & Yi, 1994, 1995; Abramowicz et al., 1995; Narayan & McClintock, 2008; Yuan & Narayan, 2014). What little radiation is present in this accretion state has hardly any effect on the dynamics of the optically thin gas. Sgr A* is a classic example of a SMBH accreting in ADAF mode (Yuan et al., 2003). Low $\,{f_{\rm Edd}}$ systems often accumulate a strong magnetic field at small radii around the black hole to form a Magnetically Arrested Disk (MAD, Igumenshchev et al., 2003; Narayan et al., 2003). In this state, they produce relativistic jets extending to thousands of gravitational radii $r_{g}=2GM_{\bullet}/c^{2}$ if the black hole is spinning (Tchekhovskoy et al., 2011). It is likely that most black holes, at least in the local Universe, accrete in the ADAF mode and are in the MAD state (Narayan et al., 2022). The accretion mode alternative to the MAD state is aptly named SANE (Standard And Normal Evolution).
@@
For $\,{f_{\rm Edd}}\gtrsim 1$, the accretion disk is radiatively inefficient because of photon trapping (Begelman, 1979; Abramowicz et al., 1988). The disk is geometrically thick, prompting the name “slim disk” for these systems. Crucially, the slim disk regime is thermally stable (Abramowicz et al., 1995; Narayan & Yi, 1995; Chen et al., 1995), which facilitates theoretical studies, especially using numerical simulations (Sądowski et al., 2014; McKinney et al., 2014). Jets are prominent if the disks reach the MAD state and if the black holes are spinning (Narayan et al., 2017; Curd & Narayan, 2023; Ricarte et al., 2023). Such systems will be extremely X-ray-bright for pole-on observers because of relativistically beamed radiation from the hot disk and jet (Curd & Narayan, 2019).
@@
Radiation plays a critical role in both the dynamics and thermodynamics of standard and slim disks (i.e., for the entire range $\,{f_{\rm Edd}}\gtrsim 0.01$). There is a strong coupling between the gas and the radiation, while radiative pressure inflates the disk in the vertical direction as $\,{f_{\rm Edd}}$ approaches or exceeds unity (Abramowicz et al., 1988). Additionally, jets and winds driven only by radiation can occur in super-Eddington systems (Sądowski et al., 2014). For these reasons, a GRRMHD code, i.e., a GRMHD code with a self-consistent treatment of the radiation field, is required. The work reported in this paper uses the GRRMHD code KORAL (Sądowski et al., 2013, 2014).
This primer has highlighted the physical differences characterizing the three fundamental accretion regimes. GRRMHD-derived spectral models are available in the ADAF regime ($\,{f_{\rm Edd}}\lesssim 0.01$) and, albeit more scarcely, in the super-Eddington regime ($\,{f_{\rm Edd}}\gtrsim 1$). However, accretion rates in between are challenging to simulate. For $\,{f_{\rm Edd}}\lesssim 0.1$, disks are geometrically very thin and require extremely high resolution, making them challenging for GRRMHD codes. On the other hand, for $\,{f_{\rm Edd}}\gtrsim 0.1$ and especially as $\,{f_{\rm Edd}}\to 1$, although disks are geometrically thicker and potentially accessible to GRRMHD codes, they are thermally unstable, as already mentioned (Shakura & Sunyaev, 1976). Only a handful of full GRRMHD simulations have been successfully run in this accretion regime; in these few simulations, strong magnetic fields stabilize the accretion flow (Sądowski, 2016) and a curious “puffy disk” structure is observed (Lančová et al. 2019; Wielgus et al. 2022). The models so far are limited to non-spinning black holes.
In the present work, we focus on the super-Eddington regime ($\,{f_{\rm Edd}}\gtrsim 1$), where the accretion disk is stable, and we have both GRRMHD codes to simulate the gas magnetohydrodynamics and radiation field of the disk (e.g., Sądowski et al., 2013, 2014) as well as post-processing codes to compute spectra (e.g., Zhu et al., 2015; Narayan et al., 2016). The results we report here are, however, valid only for the super-Eddington regime and cannot be extended to thermally unstable $\,{f_{\rm Edd}}\lesssim 1$ systems. The latter regime, though of great interest for modeling traditional quasars, is beyond the reach of our codes. Simulating them would require entirely different simulation techniques (that do not yet exist) and theoretical assumptions.
@@ 3.1 GRRMHD Simulations with KORAL
The super-Eddington accretion process onto SMBHs is simulated using the GRRMHD code KORAL (Sądowski et al., 2013, 2014), which incorporates the effects of gas dynamics, magnetic fields, and radiation within a fixed gravitational field. These simulations are performed with the Kerr metric in Kerr-Schild coordinates and employ the M1 closure method (Levermore, 1984) to handle radiative transfer, along with a radiative viscosity term to address limitations inherent in the M1 scheme (Sądowski et al., 2015). The simulations include radiative processes such as synchrotron emission, free-free and bound-free (taken from Sutherland & Dopita, 1993) emission and absorption, and Compton scattering. All these simulations are performed in 3D to correctly resolve the magnetorotational instability (Balbus & Hawley, 1991).
@@ 3.2 Radiation Post-Processing with HEROIC
In the present work, post-processing involves two stages. First, HEROIC takes the time-averaged (over the final $5000t_{g}$) and axisymmetrized data from KORAL for the gas density, temperature, and four-velocity, as well as the magnetic field strength. To be conservative, it caps the gas temperature [^2] This temperature cap prevents numerical errors in isolated cells in KORAL from causing artificially high temperatures that affect the inverse Compton scattering in the output SEDs. The calculated spectra hardly change relative to not capping the temperature, but we do it for safety. at $10^{9}$ K, which is the typical coronal temperature in Seyfert galaxies (see, e.g., Akylas & Georgantopoulos 2021). Using this information and the appropriate Kerr spacetime metric, the code iteratively solves for the frequency-dependent radiation field at each grid point, including all the emission, absorption, and scattering opacities and the effect of relativistic Comptonization. At the conclusion of this stage, the code generates a source function at each position on the grid. Using this source function, in the second stage HEROIC performs general relativistic ray-tracing, again with all opacities and scattering included, to compute the observed spectra from different viewing angles. Note that HEROIC only calculates the continuum spectrum. It does not model relativistically broadened X-ray iron lines or related reflection phenomena (e.g., Reynolds & Nowak, 2003).
The observed SEDs are re-normalized so that their bolometric luminosity matches the value expected from the Eddington ratio of each run, given a fixed black hole mass of $10^{7}\,\,{\rm M_{\odot}}$. This task required a consistent re-normalization across all runs of a factor of $10$ for the radiation field, suggesting that KORAL overestimated the temperature field by a factor of $10^{1/4}\approx 1.8$. This procedure is standard practice to correct discrepancies in calculating the radiation and temperature fields performed by KORAL, using the M1 closure method. A similar radiation and temperature field matching was also applied to other studies based on KORAL and HEROIC, e.g., Narayan et al. (2017). We also note that the fact that a consistent re-normalization was required for all runs suggests that our simulations are consistent with each other.
@@ 3.3 Description of the Simulations
The GRRMHD simulations vary the Eddington ratio, ranging from mildly super-Eddington (i.e., $\,{f_{\rm Edd}}=1.4$) to strongly super-Eddington (i.e., $\,{f_{\rm Edd}}=13.4$). Three values of the dimensionless SMBH spin parameter $a$ are considered: 0, 0.68, 0.9. By definition, $a$ is required to lie in the range $0\leq|a|\leq 1$, where $a=0$ is a Schwarzschild (i.e., non-rotating) black hole, and $a=1$ is a maximally rotating black hole (although the limiting value $a=1$ is not achievable in practice, Thorne 1974).
These 12 simulations are then post-processed to obtain the emerging spectral luminosity for 8 different viewing angles $i$ of the observer relative to the pole. The inclination angle varies between $i=10^{\circ}$ to $i=80^{\circ}$, in intervals of $10^{\circ}$. Hence, a grid of $12$ simulations $\times$ $8$ angles leads to a total of $96$ SEDs, which are used to investigate the physical parameter spaces relevant to our problem. A summary of the physical properties of the 12 simulations performed is presented in Table 1.